# Interpreting GNN-based IDS Detections Using Provenance Graph Structural Features

### v5

@@ 2.3 Explainability and Security Applications
Explaining ML-based Security Models. Due to the importance of explainability in the security domain, several explainers have been proposed different explanability framework for ML-based security analysis. LEMNA [guo2018lemna] focused on classifying PDF malware and detecting a function’s entry point in binary code using regression mixture models as a localized surrogate to approximate the classifier’s decision boundary. Recent works such as CFGExplainer [herath2022cfgexplainer] and FCGAT [someya2023fcgat] use deep surrogate models to explain GNN-based malware detection using control flow graphs and function call graphs, respectively. These methods exclusively work on homogeneous graphs, thus cannot be directly applied to heterogeneous provenance graphs.
@@
Jacob et al. [jacobs2022ai] proposed TRUSTEE, a framework that generates decision tree (DT)-based interpretations for ML models to detect shortcut learning (e.g., problem under-specification). Their success clarifying model decisions about network packets inspired us to generalize the approach to the system provenance domain. The core challenge is that the system provenance domain relies on heterogeneous multi-attributed graph datasets, which are not natively consumed by DTs. By focusing solely on graph structure and leveraging security-oriented graph structural features, ProvExplainer enables DT-based explanations in the provenance domain.
@@
Explaining GNN Models. Recent research in GNN explainers [ying2019gnnexplainer, luo2020parameterized, yuan2021explainability] has advanced in identifying key nodes, edges, or subgraphs in GNNs, and are categorized into white-box and black-box explainers. White-box methods, e.g., GNNExplainer [ying2019gnnexplainer] and PGExplainer [luo2020parameterized], access GNN internals, including model weights and gradients. Conversely, black-box methods like SubgraphX [yuan2021explainability] operate on model inputs and outputs, reducing coupling between the explanation framework and model architecture.  [kosan2023gnnx] has shown high variance in explanations provided by traditional GNN explainers, raising reliability and applicability concerns.
2.4 Ground-truth Verification
@@
Ground truth verification of ML model decisions in security-critical tasks has garnered significant attention, underscoring the role of explanations in unveiling the truth [warnecke2019don, warnecke2020evaluating, ganz2023hunting]. These studies highlight the effectiveness of white-box techniques in malware detection and vulnerability discovery, with further advancements in reconstructing ground truth around local explanations [ganz2023hunting].
@@
In system provenance, “ground truth” refers to the real-world information against which the validity of a model’s predictions are checked. In this paper, we approximate the ground truth using documentation created by security vendors, tech reports, and previous studies. The relevant processes, files, and network sockets mentioned in the documentation are designated as documented entities. We understand that the documentation provided by the security vendors can have experimental errors, as well as selection and experimental bias. To mitigate these problems, we aggregated information from multiple sources. We extract documented entities with three methods: (1) referring to pre-existing malware profile databases that contain information from different security vendors, such as VirusTotal [virustotal], we obtain activity summaries detailing network communications, file system actions, and process behaviors; (2) we extract key entities (i.e., process involved, files created and connections made) from tech reports [darpa:ground, darpa:ground2] and system manuals [linux:man]; (3) we consult dataset authors [mukherjee2023sec] and review dataset documentation to identify the components of each attack present in the datasets.
3 Problem Statement and Threat Model
@@
Our research addresses explainability in GNN-based security models built on system provenance graphs, tackling a core issue in the security domain. The complexity of explaining GNN decisions is exacerbated by graph structural learning, which adds to the inherent complexity of Neural Networks (NNs). Existing studies on GNN explainability [yuan2020explainability, gnnexplainer, subgraphx, pgexplainer, herath2022cfgexplainer, ganz2023hunting] often fail to effectively map back to system behaviors in the provenance domain. To bridge this gap, our framework employs a surrogate DT-based method that utilizes interpretable security-aware graph structural features (i.e., discriminant subgraph patterns and graph structural features) to explain GNN decisions. A graph explanation method $\mathcal{E}$ will yield a total ordering over $\mathcal{V}$ for a given graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and graph model $\mathcal{M}$.
@@
Our threat model assumes the integrity of on-device data collection, relying on provenance records secured by existing mechanisms ([provdetector, sigl2021sec, nodoze, mukherjee2023sec, liu2018ndss, cheng2024kairos, rehman2024flash, goyal2024rcaid]). Our primary objective is to generate security-aware explanations to aid security analyst and increase their trust in the GNN’s decisions. We consider graph-level anomaly detection tasks; explaining GNN decisions in node/edge level tasks is outside the scope of this work. Systematically aggregating an accurate and trustworthy ground truth for malware and APT behaviors pose a challenging open problem. In this paper, we approximate the ground truth using publicly available documentation (§ 2.4). In line with recent literature on GNN explanation [herath2022cfgexplainer, warnecke2019don, warnecke2020evaluating], adversarial samples are outside the scope of the paper. Creating robust detection and explanation systems that can withstand adversarial manipulation [goyal2023sometimes, mukherjee2023sec], procedural dataset poisoning, and detection model manipulate are critical open research problems that are orthogonal to our work.
4 ProvExplainer Overview
@@
Stage 1: Extract Security-aware Features (§ 4.1). We extract security-relevant subgraph patterns using a discriminative subgraph mining approach and graph structural features that facilitate distinction between benign and anomalous datasets. These subgraph patterns enable the identification of localized attack vectors, while global structural features capture holistic structural changes.
Stage 2: Train an Interpretable Surrogate Model (§ 4.2). Next, we utilize diverse system provenance dataset to train an interpretable surrogate DT that aligns with the GNN decisions using the extracted features. Since the GNN’s decision-making process relies on a structure-dependent message-passing algorithm, we efficiently mimic the decision-making process of the GNN’s using the structure-based features.
Stage 3: Interpret the Surrogate Model (4.3). To extract the explanation for a detection using the surrogate DT, we extract the graph nodes participating in the subgraph patterns that contribute to the DT’s decision. These explanations are valid only when the surrogate agrees with the GNN.
4.1 Feature Extraction
1:Subgraph patterns, subgraph pattern sizes, subgraph pattern support, and $K$ to select the top $K$
2:Discriminative subgraph patterns
3:Aggregate subgraphs for anomaly and benign datasets in training set
4:Calculate coverage and diversity for each pattern
5:Assign scores using for each pattern based on diversity (div), coverage (cov), pattern size (size), and pattern support (support):
$Score=\alpha\cdot\texttt{div}+\beta\cdot\texttt{cov}+\gamma\cdot\texttt{size}+\eta\cdot\texttt{support}$
6:Compute score difference for each pattern between benign and anomaly datasets
7:Select top-$K$ patterns with largest score differences using argmax_pattern
8:Save selected features and associated scores
@@ 4.1 Feature Extraction
4.1.1 Subgraph Features
The extraction of subgraph features is pivotal in distinguishing anomalous patterns from benign patterns within provenance datasets. To avoid information leakage and ensure no data snooping, patterns are mined exclusively on the training dataset, with a clear separation between training and testing data. This approach enables unbiased pattern discovery and allows the identified discriminative features to be evaluated solely on the testing dataset, maintaining a rigorous evaluation protocol.
Each file in the training dataset undergoes frequent subgraph pattern mining using a graph evolution miner (GERM [jiang2013survey]), where subgraphs of specified pattern sizes and support thresholds are identified. Support means how many of the subgraph instances were identified in the graph. As described in Algorithm 1, from each file, subgraphs are extracted based on structural pattern, generating unique patterns termed subgraph keys. These keys serve as representative features, capturing the frequency and structure of motifs present in the datasets.
@@ 4.1.1 Subgraph Features
We also measure the coverage of a subgraph pattern by measuring its presence across files within the dataset, providing a statistical representation of how commonly the pattern occurs in the dataset as a whole. Calculated as the proportion of files in which the pattern appears with non-zero support, coverage serves as an indicator of the pattern’s prevalence or rarity in benign and anomalous instances. High coverage patterns suggest commonly occurring structures within the respective dataset, representing anomalous or benign operations. We want to identify patterns that have high coverage in anomalous but low coverage in benign datasets, providing a basis for distinguishing between them.
@@
By incorporating coverage into the scoring function, we ensure that patterns contributing significantly to the anomalous dataset’s structure are prioritized, enhancing the model’s ability to generalize from typical behaviors to rare distinctive malicious behavior, indicative patterns. To quantify the discriminative utility of each pattern, we define a scoring function parameterized by weights $\alpha,\beta,\gamma,\text{ and }\eta$, which represent the relative importance of diversity, coverage, pattern size, and support, respectively. The scoring function is formulated as follows:
@@
$Score=\alpha\cdot\texttt{div}+\beta\cdot\texttt{cov}+\gamma\cdot\texttt{size}+\eta\cdot\texttt{support}$
where div denotes the calculated diversity of the pattern, cov represents the proportion of files in which the pattern appears, size is the average number of nodes in the pattern, and support reflects the frequency of occurrence of the pattern within each file. For our study, we used the following hyperparameters: $\alpha=0.2$, $\beta=0.4$, $\gamma=0.1$, and $\eta=0.3$. These values were selected to prioritize coverage and support as the primary differentiators, while also ensuring that the extracted patterns are diverse and large in size, in that order of importance.
The scoring function is applied separately to the benign and anomaly datasets in the training set. Patterns that maximize the absolute score difference between benign and anomaly datasets are selected as discriminative patterns, as they are likely to differentiate between benign and anomalous behaviors effectively. The top $k$ patterns with the largest score differences are retained and subsequently evaluated on the testing dataset, thus ensuring that the pattern discovery and evaluation are conducted independently.
4.1.2 Graph Structural Features
@@
From graph theory literature, we selected the following structural features [zaki2014data]: degree centrality, closeness centrality, betweenness centrality, eigenvector centrality, clustering coefficient, and clustering triangles. These standard features are known for capturing the key characteristics of graphs, serving as a custom signature that effectively reflects the global properties of a graph. To refresh the reader, we provide informal definitions of each feature and what is the expected trend in system graphs.
@@
Degree Centrality. Degree centrality is simply the degree of the node, or the count of its incident edges. The degree centrality of a node indicates the extent of its direct connections in the graph. When aggregated across all nodes, high degree centrality values suggest a densely connected system with significant interactivity, whereas low values indicate sparsely connected regions, potentially reflecting isolated or specialized nodes. High degree centrality could suggest the presence of core process hubs, while low values imply a more dispersed system with no central process hubs. However, degree centrality alone cannot definitively indicate malicious behavior. For example, a long-running attack creating a long process chain show low degree centrality, while attacks that interact with many files or fork multiple child malware processes will exhibit high degree centrality. In the case of a cryptominer masquerading as schtasks.exe, we would expect the benign schtasks.exe to show low centrality as it merely schedules tasks, while the cryptominer malware would have high centrality due to spawning multiple malicious processes.
@@
Closeness Centrality. Closeness centrality is the inverse of the mean shortest path distance from all other nodes to the target node, indicating how close the node is within the graph. Closeness centrality measures how centrally located a node is within the graph by examining its proximity to other nodes. High closeness centrality across the graph suggests that many nodes are relatively close to each other, indicating efficient communication pathways. In contrast, low closeness centrality values highlight dispersed sections of the graph, potentially representing decentralized or peripheral interactions within the system. High closeness centrality indicates well-connected process hubs central to the system, while low values suggest that processes are widely dispersed with limited central hubs. Empirically, system processes often have high closeness centrality, while user-level or malware processes show lower values. For instance, botnets masquerading as system-level processes will be distinguishable, since botnet will have a higher closeness centrality due to multiple network connecting being created characteristic of their DDoS activities.
@@
Betweenness Centrality. Betweenness centrality measures the frequency with which a node appears on the shortest paths between other nodes, highlighting its role as a waypoint. It indicates how often a node acts as a bridge along the shortest paths between other nodes. High betweenness centrality suggests a system’s reliance on certain core hubs or bottleneck paths, while lower values indicate multiple alternative pathways and a more distributed structure. High betweenness centrality indicate processes that serve as intermediaries in data or control flow, whereas low values suggest a network of distributed pathways with minimal reliance on specific process hubs. Malware like worms and spyware tend to have low betweenness centrality, as they seek to spread across the system and create multiple information pathways without centralizing the flow.
@@
Eigenvector Centrality. Eigenvector centrality assigns higher values to nodes connected to other influential nodes, reflecting the prominence of node clusters that are influential. High eigenvector centrality in processes indicates hubs with extensive control over information flow, while low values imply a flatter structure without prominent hubs. Ransomware tends to have high eigenvector centrality, as malicious processes involved in reading and encrypting files act as critical nodes. By keeping few child processes, ransomware conserves system resources, allowing high-volume file encryption without overloading the system.
@@
Clustering Triangles. Clustering triangles count the number of triangles that include the node. The clustering triangles count represents the frequency of triangular connections in the graph. High clustering triangle values indicate localized connectivity, suggesting shared resources or close interactions among nodes, while low values indicate a more hierarchical or linear flow. Malware often creates prominent triangles when, for example, they write template files and repeatedly execute them, making copies of themselves. Similarly, legitimate programs like svchost.exe (a service scheduler) may form triangles when creating conhost.exe and reading common *.dll files. In contrast, ransomware typically reads files to encrypt them, forming a “spoke-and-wheel” shape rather than triangles, resulting in lower triangle counts than benign processes like svchost.exe.

### v7

@@ II-C Explainability and Security Applications
Explaining ML-based Security Models. Due to the importance of explainability in the security domain, several explainability frameworks have been proposed for ML-based security solutions. LEMNA [guo2018lemna] focused on classifying PDF malware and detecting a function’s entry point in binary code using regression mixture models as a localized surrogate to approximate the classifier’s decision boundary. CFGExplainer [60] use deep surrogate models to explain GNNs but work exclusively work on homogeneous graphs.
@@
Jacob et al. [jacobs2022ai] proposed TRUSTEE, a framework that generates decision tree (DT)-based interpretations for ML models to detect shortcut learning. Their success in clarifying model decisions about network packets inspired us to generalize the approach to the system provenance domain. System provenance domain relies on heterogeneous multi-attributed graph datasets, which are not natively consumed by DTs.
@@
Explaining GNN Models. Recent research in GNN explainers [31, 32, 33] has advanced in identifying key nodes, edges, or subgraphs in GNNs, and are categorized into white-box and black-box explainers. White-box methods, e.g., GNNExplainer [31] and PGExplainer [32], access GNN internals, including model weights and gradients. Conversely, black-box methods like SubgraphX [33] operate on model inputs and outputs, reducing coupling between the explanation framework and model architecture. [62] has shown high variance in explanations provided by traditional GNN explainers, raising reliability and applicability concerns.
II-D Ground-truth Verification
@@
Ground truth verification of ML model decisions in security-critical tasks [34, 35, 36] has garnered significant attention, highlighting the role of verifying explanations in unveiling the model’s decision-making process. These studies highlight the effectiveness of white-box techniques in malware and vulnerability discovery, and reconstructing ground truth around local explanations [36].
In system provenance, “ground truth” refers to the real-world system information (e.g., system actions) against which the validity of a model’s predictions is checked. In this paper, we approximate the ground truth using documentation created by security vendors, tech reports, and previous studies. The relevant processes, files, and network sockets and their associated events mentioned in the documentation are designated as documented entities.
@@
We extract documented entities with three methods: (1) we refer to pre-existing malware profile databases that contain information from different security vendors, such as VirusTotal [30], to obtain activity summaries detailing network communications, file system actions, and process behaviors; (2) we extract key entities (i.e., process involved, files created and connections made) from tech reports [28, 29] and system manuals [63]; (3) we consult dataset authors [46] and review dataset documentation to identify the components of each attack present in the datasets. We understand that the documentation provided by the security vendors can have experimental errors, as well as selection and experimental bias. To mitigate these problems, we aggregated information from multiple sources.
III Motivation
In this section, we present a real-world attack scenario from the DAPRA TC [29] dataset to highlight the limitations of existing GNN explainers and demonstrate the effectiveness and utility of ProvExplainer. The experimental setup to evaluate ProvExplainer is described in § VI and more case studies are described in § VII and appendix (§ A-E and § A-F).
(a) Attack Graph.
(b) SOTA GNN explainations.
(c) Pattern #1 and Pattern #2.
Attack Scenario (Trace: Phishing Email [darpa]). Figure 1 presents a simplified depiction of the APT attack initialized by phishing e-mail link. This particular attack was executed as part of a DAPRA red team exercise by Trace team. First, the attacker launches a phishing campaign that compromises the identity of an employee. Leveraging this compromised identity, the attacker then sends deceptive emails containing links to a malicious website (i.e., www.nasa.ng) to other employees. When an employee visits the website, a Trojan is downloaded and placed in /tmp/vUgefal. Upon execution, the Trojan replicates itself within the system by reading from its own file.
The malware adopts the name firefox nasa.ng and stores sensitive content in /usr/share/locale/locale.alias for exfiltration. To evade detection, the Trojan creates additional malicious copies using benign program names (e.g., firefox), allowing it to replicate unhindered and overwhelm the system with malicious processes. These malware clones quietly read sensitive user files while the original Trojan remains undetected.
Limitation of SOTA GNN explainers. 1(b) compares the explanations produced by ProvExplainer with those of SOTA GNN explainers (i.e., GNNExplainer, PGExplainer, and SubgraphX) in capturing malicious activities relevant to GNN-based IDS. While GNNExplainer effectively uncovers the initial compromise behavior such as payload dropping and staging, it fails to capture later attack stages notable the interactions with shared libraries and cache files. PGExplainer, in contrast, highlights irrelevant benign system components like xfce4-appfinder (invoked by a DARPA script to simulate an enterprise environment) but overlooks critical malicious interactions. SubgraphX accurately detects the malicious inheritance behavior of Firefox creating multiple malware clones, yet it misses key points such as how the malware writes its payload to disk and accesses sensitive files.
ProvExplainer presents a complete view of the kill chain from Trojan deployment and clone creation to the final step of accessing sensitive data. ProvExplainer uses security relevant discriminative subgraph patterns and graph structural metrics as features to capture the holistic malicious behavior within the system. Moreover, it accomplishes this despite functioning as a black-box solution, requiring neither direct access to the underlying model nor the training data. This flexibility makes ProvExplainer compatible with a wide range of GNN-based IDS deployments.
ProvExplainer Features: Clustering coefficient (graph structural metric); pattern #1 and #2 (discriminative subgraph patterns).
System Interpretation of ProvExplainer Features. In this APT scenario, clustering coefficient, pattern #1 and pattern #2 in 1(c) showcase the malware’s structures. Pattern #1 reveals the Trojan’s propagating activity, where initial malware (Process P1) reads the firefox binary (/usr/local/bin/firefox) to execute it and write the maliclious payload (/tmp/vUgefal) to a file that is read by other malicious firefox processes (Process P4, P7,P8, P9). A series of intra-malware communication happens when the another malicious firefox (Process P2) reads the malicious payload and writes to system file ( /etc/nsswitch.conf). As seen in pattern #2, other malware processes (Process P4, P5, P6) read the system file to update their configurations and task lists to better attack the victim. This propagation of information continues such that any subsequent malware processes that read from the system file gets the updated configuration.
By masquerading as benign programs, the malware disguises its malicious processes, allowing it to proliferate undetected. The clustering coefficient of the graph increases because of multiple malware reading the template files by forming large process creation and file read clusters.
IV Problem Statement and Threat Model
@@
Our research addresses explainability in GNN-based security models built on system provenance graphs, tackling a core issue in the security domain. The complexity of explaining GNN decisions is exacerbated by graph structural learning, which adds to the inherent complexity of Neural Networks (NNs). Existing studies on GNN explainability [43, gnnexplainer, subgraphx, pgexplainer, 60, 36] often fail to effectively map back to system behaviors in the provenance domain. To bridge this gap, ProvExplainer, employs a surrogate DT-based method that utilizes interpretable security-aware graph structural features (i.e., discriminant subgraph patterns and graph structural features) to explain GNN decisions. Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and a GNN model $\mathcal{M}$, an explanation method $\mathcal{EM}$ produces:
  • Set of subgraphs $\mathcal{SG}$, where $\mathcal{SG}=\{\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{k}\}$ and $\mathcal{G}_{i}=(\mathcal{V}_{i},\mathcal{E}_{i})$ for $i=1..k$, with $\mathcal{V}_{i}\subseteq\mathcal{V}$ and $\mathcal{E}_{i}\subseteq\mathcal{E}$.
  • Set of properties $\mathcal{P}rop=\{p_{1},p_{2},\dots,p_{j}\}$, where each $p_{j}$ is a graph structural property of $\mathcal{G}$.
The subgraphs $\mathcal{SG}$ and properties $\mathcal{P}rop$ are designed to highlight the most relevant components of $\mathcal{G}$ that contribute to the predictions of $\mathcal{M}$.
@@
Our threat model assumes the integrity of on-device data collection, relying on provenance records secured by existing mechanisms ([provdetector, 18, nodoze, 46, 65, 19, 20, 21]). Our primary objective is to generate security-aware explanations to aid security analyst and increase their trust in the GNN’s decisions. We consider graph-level anomaly detection tasks; explaining GNN decisions in node/edge level tasks is outside the scope of this work. Systematically aggregating an accurate and trustworthy ground truth for malware and APT behaviors poses a challenging open problem. In this paper, we approximate the ground truth using publicly available documentation (§ II-D). In line with recent literature on GNN explanation [60, 34, 35], adversarial samples are outside the scope of the paper. Creating robust detection and explanation systems that can withstand adversarial manipulation [goyal2023sometimes, 46], dataset poisoning, and detection model manipulations are critical open research problems that are orthogonal to our work.
V ProvExplainer
@@
Stage 1: Extract Security-aware Features (§ V-A). We extract security-relevant features such as subgraph patterns using a discriminative subgraph miner [38] and graph structural features such as clustering and centrality measures that facilitate distinction between benign and anomalous datasets. The subgraph patterns enable the identification of localized attack vectors, while global graph structural features capture holistic attack behaviors.
Stage 2: Train an Interpretable Surrogate Model (§ V-B). Next, we utilize diverse system provenance dataset to train an interpretable surrogate DT that aligns with the GNN decisions using the extracted features. Since the GNN’s decision-making process relies on a structure-dependent message-passing algorithm, we efficiently mimic the decision-making process of the GNN’s using the security-relevant features.
Stage 3: Interpret the Surrogate Model (V-C). To interpret the explanation for detection using the surrogate DT, we extract the graph nodes participating in the subgraph patterns that contribute to the DT’s decision. These explanations are valid only when the surrogate agrees with the GNN.
V-A Feature Extraction
@@
1:Benign $D_{b}$ and Anomaly $D_{a}$ datasets, $k$, and score_funcscore\_func
2:Top-$k$ Discriminative Subgraph Patterns
[…]
5:for each pattern $p\in CP$ do
6:  $div\leftarrow\text{getDIVERSITY}(p)$
7:  for each graph $G\in D_{b}$ do
8:    $size_{b},support_{b}\leftarrow\text{getSIZEandSUPPORT}(p,G)$
9:  end for
10:  $cov_{b}\leftarrow\text{getCOVERAGE}(p,D_{b})$
11:  $score_{b}\leftarrow score\_func(div,cov_{b},size_{b},support_{b})$
12:  for each graph $G\in D_{a}$ do
13:    $size_{a},support_{a}\leftarrow\text{getSIZEandSUPPORT}(p,G)$
14:  end for
15:  $cov_{a}\leftarrow\text{getCOVERAGE}(p,D_{a})$
16:  $score_{a}\leftarrow score\_func(div,cov_{a},size_{a},support_{a})$
17:  $Patterns.\text{append}([p,|\text{score}_{a}-\text{score}_{b}|])$
18:end for
19:Sort PatternsPatterns in desc order based on the score difference
20:return $Patterns[:k]$
Security-relevant Features. The extraction of security-relevant subgraph features is pivotal in distinguishing anomalous patterns from benign patterns within provenance datasets. To avoid information leakage and ensure no data snooping, patterns that are mined exclusively on the training dataset are used for explanation evaluation. This approach enables unbiased pattern discovery and allows the identified discriminative features to be evaluated solely on the testing dataset, maintaining a rigorous evaluation protocol.
Frequent subgraph patterns are mined from the training dataset using a graph evolution miner (GERM [39]), where subgraphs of specified sizes and their supports are identified. Support means how many times the subgraph instances were identified in the graph. These patterns serve as representative features, capturing the frequency and structure of motifs present in the datasets.
@@
We also measure the coverage of a subgraph pattern by measuring its presence across graphs within the dataset, providing a statistical representation of how commonly the pattern occurs in the dataset as a whole. Calculated as the proportion of graphs in which the pattern appears with non-zero support, coverage serves as an indicator of the pattern’s prevalence or rarity in benign and anomalous instances. High coverage patterns suggest commonly occurring structures within the respective dataset, representing anomalous or benign operations. We want to identify patterns that have high coverage in anomalous but low coverage in benign datasets, providing a basis for distinguishing between them.
@@
To quantify the discriminative utility of each pattern, we define a scoring function parameterized by weights $\alpha,\beta,\gamma,\text{ and }\eta$, which represent the relative importance of diversity, coverage, pattern size, and support, respectively. The score_func in Algorithm 1 (i.e., scoring function) is:
@@
$\text{Score}=\alpha\cdot\texttt{div}+\beta\cdot\texttt{cov}+\gamma\cdot\texttt{size}+\eta\cdot\texttt{support}$
where div denotes the diversity of the pattern, cov represents the proportion of graphs in which the pattern appears, size is the average number of nodes in the pattern, and support reflects the frequency of pattern occurrence within each graph.
Increasing $\alpha$ enhances the selection of patterns with diverse node types, which may improve the ability to identify unique patterns but could reduce the focus on more frequent patterns. Raising $\beta$ results in higher scores for patterns that appear across a larger number of graphs, promoting generalization but potentially overlooking rare but important patterns. Elevating $\gamma$ favors patterns with larger node counts, which can capture more complex interactions but might exclude simpler yet crucial substructures. Finally, increasing $\eta$ prioritizes patterns with higher frequency, amplifying commonly occurring patterns but at the expense of pattern diversity.
For our study, we used the following weights: $\alpha=0.2$, $\beta=0.4$, $\gamma=0.1$, and $\eta=0.3$. These values were selected empirically to prioritize coverage and support as the primary differentiators, while also ensuring that the extracted patterns are diverse and large. Subgraph patterns are extracted from each graph in the dataset using Algorithm 1 but the discriminant subgraph patterns are selected exclusively from the anomalous dataset, as we ensure that patterns contributing significantly to the anomalous dataset’s structure are prioritized. Patterns that maximize the score difference between benign and anomaly datasets are selected as discriminative patterns, as they effectively differentiate between benign and anomalous behaviors. Also, subgraph patterns are extracted solely from the training set and evaluated on an independent test set, thereby eliminating any risk of data leakage.
@@
Graph Structural Features. From graph theory literature, we selected the following structural features [67]: degree centrality, closeness centrality, betweenness centrality, eigenvector centrality, clustering coefficient, and clustering triangles. These standard features are known for capturing the key characteristics of graphs, serving as a custom signature that effectively reflects the global properties of a graph. To refresh the reader, we provide informal definitions and their meaning in system provenance domain.
@@
Degree Centrality. Degree centrality is simply the degree of the node, or the count of its incident edges. The degree centrality of a node indicates the extent of its direct connections in the graph. When aggregated across all nodes, high degree centrality values suggest a densely connected system with significant interactivity, whereas low values indicate sparsely connected regions, potentially reflecting isolated or specialized nodes. High degree centrality could suggest the presence of core process hubs, while low values imply a more dispersed system with no central process hubs. However, degree centrality alone cannot definitively indicate malicious behavior. For example, a long-running attack creating a long process chain shows a low degree of centrality, while attacks that interact with many files or fork multiple child malware processes will exhibit a high degree of centrality. In the case of a cryptominer masquerading as schtasks.exe, we expect the benign schtasks.exe to show low centrality as it schedules tasks, while the cryptominer malware have high centrality due to spawning multiple malicious processes.
@@
Closeness Centrality. Closeness centrality is the inverse of the mean shortest path distance from all other nodes to the target node, indicating how close the nodes are within the graph. Closeness centrality measures how centrally located a node is within the graph by examining its proximity to other nodes. High closeness centrality across the graph suggests that many nodes are relatively close to each other, indicating efficient communication pathways. In contrast, low closeness centrality values highlight dispersed sections of the graph, potentially representing decentralized or peripheral interactions within the system. High closeness centrality indicates well-connected process hubs central to the system, while low values suggest that processes are widely dispersed with limited central hubs. Empirically, system processes often have high closeness centrality, while user-level or malware processes show lower values. For instance, botnets masquerading as system-level processes will be distinguishable, since botnet will have a higher closeness centrality due to multiple network connecting being created due to their DDoS activities.
@@
Betweenness Centrality. Betweenness centrality measures the frequency with which a node appears on the shortest paths between other nodes, highlighting its role as a waypoint. It indicates how often a node acts as a bridge along the shortest paths between other nodes. High betweenness centrality suggests a system’s reliance on certain core hubs or bottleneck paths, while lower values indicate multiple alternative pathways and a more distributed structure. High betweenness centrality indicate processes that serve as intermediaries in data or control flow, whereas low values suggest a network of distributed pathways with minimal reliance on specific process hubs. Malware like worms and spyware tend to have low betweenness centrality, as they seek to spread across the system and create multiple distributed information pathways.
@@
Eigenvector Centrality. Eigenvector centrality assigns higher values to nodes connected to other influential nodes, reflecting the prominence of node clusters that are influential. High eigenvector centrality in processes indicates hubs with extensive control over information flow, while low values imply a flatter structure without prominent hubs. Ransomware have high eigenvector centrality, as malicious processes involved in reading and encrypting files act as critical nodes. They keep few child processes to allow high-volume file encryption without overloading the system.
@@
Clustering Triangles. Clustering triangles count the number of triangles that include the node. The clustering triangles count represents the frequency of triangular connections in the graph. High clustering triangle values indicate localized connectivity, suggesting shared resources or close interactions among nodes, while low values indicate a more hierarchical or linear flow. Malware often creates prominent triangles when they write template files and repeatedly execute them, making copies of themselves. Similarly, legitimate programs like svchost.exe (a service scheduler) form triangles when creating conhost.exe and reading standard *.dll files. In contrast, ransomware reads files to encrypt them, creating “spoke-and-wheel” shape rather than triangles, resulting in lower triangle counts than benign processes, svchost.exe.